### v2

# Speculative Safety-Aware Decoding

Abstract
Despite extensive efforts to align Large Language Models (LLMs) with human values and safety rules, jailbreak attacks that exploit certain vulnerabilities continuously emerge, highlighting the need to strengthen existing LLMs with additional safety properties to defend against these attacks. However, tuning large models has become increasingly resource-intensive and may have difficulty ensuring consistent performance. We introduce Speculative Safety-Aware Decoding (SSD), a lightweight decoding-time approach that equips LLMs with the desired safety property while accelerating inference. We assume that there exists a small language model that possesses this desired property. SSD integrates speculative sampling during decoding and leverages the match ratio between the small and composite models to quantify jailbreak risks. This enables SSD to dynamically switch between decoding schemes to prioritize utility or safety, to handle the challenge of different model capacities. The output token is then sampled from a new distribution that combines the distributions of the original and the small models. Experimental results show that SSD successfully equips the large model with the desired safety property, and also allows the model to remain helpful to benign queries. Furthermore, SSD accelerates the inference time, thanks to the speculative sampling design.
Speculative Safety-Aware Decoding
WARNING: This paper contains model outputs that may be considered offensive.
Xuekang Wang2   Shengyu Zhu1thanks: Corresponding author: zhushengyu@ict.ac.cn   Xueqi Cheng1 1 State Key Laboratory of AI Safety, Institute of Computing Technology, CAS 2 School of Computer Science & Technology, Beijing Institute of Technology wangxk@bit.edu.cn  {zhushengyu,cxq}@ict.ac.cn
1 Introduction
Large language models (LLMs) have demonstrated remarkable performances in a wide range of natural language tasks (Achiam et al., 2023; Touvron et al., 2023; Chiang et al., 2023; Team, 2023). Currently, their safety hinges on various alignment approaches (Leike et al., 2018; Kenton et al., 2021; Ji et al., 2023), including supervised fine-tuning (Wei et al., 2021) and preference-based methods (Ouyang et al., 2022; Bai et al., 2022; Rafailov et al., 2023) as commonly-used practices. These approaches optimize models to align with human values and refuse unsafe responses, to ensure LLMs are helpful and harmless. However, LLMs are still vulnerable to jailbreak attacks, which can bypass the alignment constraint and result in harmful responses (Wei et al., 2023b; Deng et al., 2023a; Li et al., 2023; Zeng et al., 2024; Zou et al., 2023; Liu et al., 2023; Yu et al., 2023; Chao et al., 2023).
Recently, many defense methods have been proposed to reduce the likelihood of generating harmful responses, by adjusting output probabilities of the aligned LLM at decoding time (Xu et al., 2024; Banerjee et al., 2025; Djuhera et al., 2025; Liu et al., 2024b). For example, SafeDecoding (Xu et al., 2024) amplifies the probability of safety token sequences and attenuates the harmful ones at the first few output tokens in a model arithmetic manner (Liu et al., 2021; Dekoninck et al., 2024; Liu et al., 2024a). While being effective in defending against jailbreak attacks, most aligned LLMs, even with these defense methods, still lack the “deep safety alignment” guarantee, an important safety property recently proposed by Qi et al. (2025). However, fine-tuning large models with augmented datasets have become increasingly resource-intensive, and may have difficulty ensuring consistent performance (Qi et al., 2023; Yao et al., 2023). As such, we ask: can we develop a lightweight and efficient approach to further strengthen the safety of LLMs with an additional, desired safety property?
Figure 1: Illustration of our approach. SSD integrates speculative sampling during decoding and dynamically switches between two decoding schemes: $\mathrm{Intersection}$ for utility and $\mathrm{Union}$ for safety. In this example with a harmful input query, the large model may respond affirmatively while the small expert model tends to output refusals. This difference would lead to a low match ratio and SSD switches to the $\mathrm{Union}$ scheme to prioritize safety.
This paper introduces a lightweight decoding-time approach that can also accelerate inference of the LLM, without incurring the cost of tuning the large model’s parameters. We assume that there exists a smaller language model that possesses the desired safety property, which can be obtained from fine-tuning or other alignment approaches. In particular, we will focus on the deep safety alignment (Qi et al., 2025) property that enables the model to stop generating harmful responses even if the safety alignment is bypassed initially.
Existing decoding-time defense methods such as Xu et al. (2024); Banerjee et al. (2025) also employ auxiliary models to improve safety, but they generally require fine-tuning models of similar sizes and need at least one inference of the LLM per output token. In contrast, a smaller model allows for a faster inference than the original model in a speculative manner (Stern et al., 2018; Chen et al., 2023; Leviathan et al., 2023). In Liu et al. (2021, 2024a); Mitchell et al. (2024), smaller models are also employed to achieve efficient fine-tuning or controllable text generation using model arithmetic. Their tasks do not directly target safety and are different from ours. Indeed, as shown in Section˜4.1, a direct implementation of model arithmetic leads to over-refusal behavior and degrades the model’s helpfulness, largely due to the inherent different capacities of the large model and the small draft model. Devising a decoding strategy in light of the difference of model capabilities would be to key to improving safety while maintaining utility.
In this work, we propose Speculative Safety-Aware Decoding (SSD) that integrates speculative sampling during decoding and leverages the match ratio between the expert and the composite models to dynamically switch between utility and safety decoding schemes, as illustrated in Figure˜1. The match ratio is defined as the agreement rate between the two models over generated tokens and is used as a way of quantifying jailbreak risks. For benign queries, both the expert and original models are likely to respond positively and the match ratio keeps high. In this case, SSD creates a sample space by taking the intersection of the top tokens from both the original and expert models. Conversely, for harmful queries, the match ratio would be low, and we enforce the additional safety property by identifying the union of the top tokens of the two models as sample space. Finally, SSD defines a new sampling distribution over the constructed sample space and then samples tokens to generate responses to user’s query.
In the experiments, we evaluate the effectiveness, helpfulness, and efficiency of the proposed method. The results show that SSD successfully equips the large model with the desired deep safety alignment property, and also allows the model to remain helpful to queries from benign users. Interestingly, for less secure models like Vicuna (Chiang et al., 2023), SSD can achieve a better safety alignment performance than directly fine-tuning the original model with supervised alignment objective. Furthermore, SSD accelerates the inference time, thanks to the speculative sampling design.
2 Related Work
Jailbreak Attacks.
Jailbreak attacks seek to bypass the safety alignment mechanisms of LLMs to elicit unsafe and harmful responses. Early attempts usually rely on manually crafted adversarial prompts that exploit the vulnerabilities of competing objective and mismatched generalization, e.g., Mowshowitz (2022); Li et al. (2023); Wei et al. (2023a); Deng et al. (2023b). In addition, Zeng et al. (2024) apply a persuasion taxonomy from social science to manipulate model responses. More recent jailbreak methods focus on optimization based methods to automate prompt generation, such as gradient based methods like Zou et al. (2023) and genetic algorithm based methods like Liu et al. (2023) a Some other approaches incorporate red-teaming strategies, by using auxiliary LLMs to assist generating and refining jailbreaks, e.g., GPTFuzzer (Yu et al., 2023) and PAIR (Chao et al., 2023). These jailbreak attacks underscore the importance of LLM safety to responsible outputs.
Jailbreak Defenses.
Many defense methods have been proposed to mitigate the above jailbreak attacks. A class of methods are to detect the harmfulness in the input queries or output responses, e.g., using keyword matching (Deng et al., 2023a), perplexity based metrics (Alon and Kamfonas, 2023; Jain et al., 2023), or a judge LLM (Helbling et al., 2023). SmoothLLM (Robey et al., 2023) identifies adversarial inputs based on multiple perturbed input copies, and RA-LLM (Cao et al., 2024) uses a robustly-aligned LLM for alignment check. EEG-Defender (Zhao et al., 2024a) trains a classifier and takes the embeddings at different layers to evaluate the harmfulness of the input query. Another class of methods reduces the likelihood of generating harmful responses. Some methods in this class involve input modification, like paraphrasing and retokenization Jain et al. (2023). Other approaches utilize prompts with question-and-answer interactions (Wei et al., 2023b; Zhang et al., 2024b) or incorporate self-reminders in system prompts to enhance responsible responses (Wu et al., 2023).
Closely related to the present work is decoding-time defense within the second class. Several works directly adjust the decoding probabilities to formulate safer outputs (Xu et al., 2024; Banerjee et al., 2025; Djuhera et al., 2025; Liu et al., 2024b; Zhao et al., 2024c). For example, Xu et al. (2024); Banerjee et al. (2025) employ fine-tuned models to assist reducing harmful responses in a model arithmetic manner. Liu et al. (2024b) define competitive index to quantify alignment failures and utilize feedback to compute new logits. These methods mostly require fine-tuning the original models and need at least one inference of the LLM per output token. Also related is Zeng et al. (2025) that trains neural networks to provide a real-time safety assessment of candidate tokens and then adjusts the top-$k$ tokens to safer alternatives. Noticeably, the above methods aim to improve the overall defense performance across various types of attacks, while our goal is to strengthen an existing LLM with a specific safety property in a lightweight and efficient way.
Controllable Generation.
Our work is also related to controllable generation that aims to introduce certain attributes in LLM outputs, which can be achieved by modifying the output probabilities to bias towards the desired attribute. In these scenarios, a strength parameter is often used to control the degree of the conditioning. Existing methods include Deng and Raffel (2023); Liu et al. (2024a, 2021); Kim et al. (2023); Pei et al. (2023), among others. These tasks commonly involve non-toxicity and positive sentiment, and their goal is different from ours: improving model safety while maintaining sufficient utility. Indeed, as we show in Section˜4.1, a direct implementation does not handle well safety and utility simultaneously.
3 Background and Problem Setting
In this section, we introduce related background and then describe our problem setup.
3.1 Background
Notation.
We denote the original auto-regressive LLM by $M$ and a smaller draft model by $m$. Let $\bm{x}_{1:n-1}$ represent a sequence of generated tokens and $x_{n}$ denote the $n$-th token. Given tokens $\bm{x}_{1:n-1}$, the sampling or decoding distribution of $M$ is represented by $P_{M}(x|\bm{x}_{1:n-1})$, which can be used to generate $x_{n}$ through various decoding strategies.
Shallow and Deep Safety Alignment.
Recently, several works Qi et al. (2025); Lin et al. (2024); Zhang and Wu (2024); Zhao et al. (2024b); Zhou et al. (2023) have revealed a critical limitation on existing safety alignment: the aligned model primarily relies on the first few output tokens, such as “I cannot” and “I apologize”, to refuse harmful queries. If the initial output tokens deviate from these safety prefixes, e.g., by prefilling attack starting with “Sure”, the model is likely to continue generating harmful responses to user’s request. This superficial alignment is referred to as shallow safety alignment in Qi et al. (2025). By deepening safety alignment, we hope that the model can recover from harmful starting conditions.
SafeDecoding.
SafeDecoding (Xu et al., 2024) is a decoding-time method that guides the original model $M$ towards generating safer outputs. It begins by fine-tuning $M$ to obtain an expert model $M^{\prime}$ with hardened safety. Then the output probabilities of both models are employed to construct new sampling distributions to reduce harmful outputs.
Specifically, let $\mathcal{V}_{n}$ and $\mathcal{V}_{n}^{\prime}$ denote the sets of tokens sampled from the original model $M$ and the expert model $M^{\prime}$ at the $n$-th decoding step, respectively. The tokens in each set are assumed to be sorted in descending order of their probabilities. A target sample space $\hat{\mathcal{V}}_{n}(c)$ is constructed as the intersection of the top $k$ tokens from $\mathcal{V}_{n}$ and $\mathcal{V}_{n}^{\prime}$:
$\hat{\mathcal{V}}_{n}(c)=\operatorname*{arg\,min}_{\mathcal{V}=\mathcal{V}_{n}(k)\cap\mathcal{V}_{n}^{\prime}(k)}k\quad\text{s.t.}\quad|\mathcal{V}|\geq c,$ (1)
where $c$ is a tunable parameter that controls the size of the sample space, and $\mathcal{V}_{n}(k)$ and $\mathcal{V}^{\prime}_{n}(k)$ are the top $k$ tokens from $\mathcal{V}_{n}$ and $\mathcal{V}^{\prime}_{n}$, respectively. As discussed in Xu et al. (2024), taking the intersection can leverage the advantages of both models. To generate the $n$-th token, the final probability function $F_{n}$ over $\hat{\mathcal{V}}_{n}(c)$ is
$F_{n}(x)$ $=P_{M}(x|\bm{x}_{1:n-1})$
$\hskip-15.0pt+\alpha\left(P_{M^{\prime}}(x|\bm{x}_{1:n-1})-P_{M}(x|\bm{x}_{1:n-1})\right),$ (2)
where $\alpha\geq 0$ is a hyperparameter. The final sampling distribution is constructed by normalizing the values in Eq. (3.1), e.g., applying Softmax to $F_{n}(x)$. For computation and generation quality concerns, this SafeDecoding procedure is applied only at the initial few output tokens in Xu et al. (2024).
Speculative Sampling.
To speed up inference of an LLM $M$, speculative sampling (Stern et al., 2018; Chen et al., 2023; Leviathan et al., 2023) employs a small, fast model to first predict several tokens $\tilde{x}_{1},\tilde{x}_{2},\cdots,\tilde{x}_{T}$, which are then verified by the large model $M$. The verification can be done in parallel and are significantly cheaper than calling the target model $T$ times due to the auto-regressive structure of LLMs. Then, if $\tilde{x}_{t}$ is rejected by the large model, the subsequent draft tokens $\tilde{x}_{i+t},\cdots,\tilde{x}_{T}$ are discarded and the output tokens would be re-sampled. This procedure can produce the same prediction of $M$ with certain decoding strategies, and achieve a faster inference if a large portion of draft tokens are accepted.
3.2 Problem Setting
In this work, we aim to strengthen a pretrained large model $M$ with an additional safety alignment property, without incurring the cost of tuning its parameters. Besides steering the model to be safer, the developed decoding strategy shall also be: 1) helpful: the resulting model outputs should maintain helpful to benign queries; 2) efficient: the approach should be both computation- and time-efficient at inference time; and 3) compatible: existing LLMs have diverse architectures and the decoding strategy shall work with different LLMs.
Our setting represents scenarios when there is a need to equip existing LLMs with new safety alignment properties in a lightweight and efficient way. In this paper, we focus on the deep safety alignment property (Qi et al., 2025), and assume that there is a small draft model $m$ that has been trained to have the desired deep alignment property, e.g., by fine-tuning the small model using an augmented harmful dataset as in Qi et al. (2025). The small model does not need to be in the same model family of $M$, but is required to share the same vocabulary. Instead of fine-tuning $M$ that may require high training resource, we would like to modify the output responses at decoding time, and at the same time, improve the inference efficiency.
4 Method
This section presents our approach that adaptively adjusts the LLM probabilities at decoding time. We first introduce our motivation and key insight, and then formally describe the proposed approach.
4.1 Motivation
Inspired by decoding-time defense methods, we utilize a small expert model $m$ to steer the large model $M$ to generate safer responses that adhere to the additional safety property, i.e., deep safety alignment in this paper. Meanwhile, the small model can act as a draft model in speculative sampling, providing a way of accelerating decoding.
A direct approach is to replacing the fine-tuned model $M^{\prime}$ in SafeDecoding with this small model $m$, and then apply speculative sampling during inference. We will refer to the resulting model as composite model in this paper. However, increasing the strength parameter $\alpha$ in Eq. (3.1) severely degrades the utility performance, while a small $\alpha$ is insufficient to equip $M$ with the desired deep alignment property.
Concretely, we use a fine-tuned TinyLlama-1.1B-Chat model (Zhang et al., 2024a) as the small model for Llama2-13b-chat Touvron et al. (2023). We test utility performance on GSM8K (Cobbe et al., 2021) and safety performance using prefilling attack of 20 tokens on the Harmful HEx-PHI data Qi et al. (2025) (detailed setup can be found in Section˜5.1). Figure˜2 validates that this direct approach cannot handle well both safety and utility. Unlike existing decoding-time methods (Xu et al., 2024; Banerjee et al., 2025; Djuhera et al., 2025), the inherent difference of model capacities places a key challenge here and we cannot rely on a single decoding scheme. Is there a way to adaptively switch decoding schemes to balance safety and utility?
Figure 2: Tradeoff between utility and safety. Higher score indicates better performance for both metrics.
4.2 Key Insight: Match Ratio
We notice that the original LLM $M$ has been trained on a vast corpus and is generally more capable, but is also more vulnerable to jailbreak attacks utilizing the shallow alignment shortcut. In contrast, the expert model $m$ is robust to such jailbreak attacks as it has been trained to possess deeper safety alignment. Consequently, when facing these attacks, $M$ is more likely to respond affirmatively while $m$ is expected to decline the response. For benign queries, both models are likely to behave positively. The difference in the decoding distributions between the two models hence provides a way of quantifying jailbreak risks.
In this paper, we use match ratio of the expert and composite models as our metric of different decoding schemes. Formally, assume output tokens $\{x_{n}\}_{n}$ that are generated by $M$ and $m$ using speculative sampling and model arithmetic (like in SafeDecoding). Denote by $I(n)\in\{0,1\}$ the indicator function of whether $x_{n}$ is drafted by $m$ and is also accepted by the composite distribution. We divide the decoded tokens into consecutive bins of size $b$. Define the match ratio of the $i$-th bin as
$\beta_{i}=\frac{1}{b}\sum_{n\in[(i-1)b+1,ib]}I(n).$ (3)
Intuitively, $\beta_{i}$ captures the agreement rate between the two models over each $b$ tokens and reflects how different the two models behave to an input query.
Figure˜3 depicts the average match ratios between TinyLlama and Llama2-13b at different bins of size $b=7$, again using GSM8K and the Harmful HEX-PHI datasets. The match ratio is clearly different between benign and harmful queries, particularly at the initial decoding phase. This validates our use of match ratio as a proper indicator of switching between different decoding schemes.
Figure 3: Mean match ratio of utility and safety tasks.
4.3 Speculative Safety-Aware Decoding
Based on the above insight, we now present our decoding strategy to handle the key challenge resulting from the different model capacities. In particular, the match ratio indicates different decoding schemes of prioritizing either utility or safety.
Utility.
While the expert and original models may respond positively to benign queries and the match ratio is high, the limited capacity of the small draft model may still degrade the model performance, as shown in Figure˜2. A reason is that the intersection operation in Eq. (1) may discard utility tokens generated by $M$. In other words, the tokens with large probabilities generated by $M$ are not necessarily among the top $k$ tokens of the small model $m$.
To mitigate the utility degradation caused by $m$, we can directly rely on $M$ when the top few tokens of $\mathcal{V}_{n}$ do not appear in the intersection $\hat{\mathcal{V}}_{n}(c)$. The new intersection set is then defined as
[…]
where $\kappa\ll c$ is a small integer compared with the size $c$ of target sample space, $\mathcal{V}_{n}(\kappa)$ is the top $\kappa$ tokens generated by $M$, and $\hat{\mathcal{V}}_{n}(c)$ is the intersection set defined in Eq. (1).
Safety.
For harmful queries, we would like to bias more towards to the expert model. However, if the safety related tokens have low probabilities under $M$, then these tokens would be discarded, again due to the intersection operation. Since we aim to introduce the additional safety property, the composite distribution should represent the union of the characteristics of both models, and here we directly take the union of the two sets as our target sample space. Specifically, for a predefined size $c$ of target sample space, we first take the top $c$ tokens from $\mathcal{V}_{n}$ and $\mathcal{V}^{\prime}_{n}$, and then compute the union as
${\mathcal{U}}_{n}=\mathcal{V}_{n}(c)\cup\mathcal{V}_{n}^{\prime}(c).$ (5)
Speculative Safety-Aware Decoding.
Our final algorithm, Speculative Safety-Aware Decoding (SSD), integrates speculative sampling with running match ratio as a measure of switching between two decoding schemes: $\mathrm{Intersection}$ and $\mathrm{Union}$, as outlined in Algorithm˜1.
SSD consists of three main stages in each loop. First, we employ the small expert model $m$ to generate $T$ draft tokens in an auto-regressive manner, and then run in parallel the large model $M$ to obtain $T$ sets of decoding probabilities, each conditioned on the incremental prefixes of the draft token sequences. Next, SSD constructs the target sample space and probability functions according to the present decoding scheme, which is either $\mathrm{Intersection}$ or $\mathrm{Union}$. An output token is then sampled according to the composite distribution. Lastly, for every $b$ output tokens, SSD computes the match ratio and decides whether to switch the decoding scheme according to a predefined threshold. We also take into account that the two models behave more similarly when conditioned on more output tokens, as seen from Figure˜3. As such, SSD adjusts the threshold and strength parameters in an annealing way if the decoding scheme keeps unchanged. Due to space limit, a detailed description is given in Algorithm˜2 in Appendix˜A.
Algorithm 1 Speculative Safety-Aware Decoding
0: original and expert models $M,m$, lookahead $T$, minimum output sequence length NN, bin size $b$, strength weights $\alpha_{I}^{0},\alpha_{U}^{0}$
1: Initialize: $\mathrm{Intersection}$, $\alpha_{I}\leftarrow\alpha_{I}^{0}$, $\alpha_{U}\leftarrow\alpha_{U}^{0}$
2:while $n<N$ do
3:  for $t=1\text{ to }T$ do
4:   Sample draft tokens auto-regressively $\tilde{x}_{t}\sim P_{m}(x|\bm{x}_{1:n},\tilde{x}_{1},\cdots,\tilde{x}_{t-1})$.
5:  end for
6:  Compute in parallel $p_{M}(x|\bm{x}_{1:n}),\cdots,$ $p_{M}(x|\bm{x}_{1:n},\tilde{x}_{1},\cdots,\tilde{x}_{T-1})$.
7:  for $t=1\text{ to }T$ do
8:   if $\mathrm{Intersection}$ then
9:    Compute sample space ${\mathcal{S}}_{n}$ as in Eq. (4), and sample $x_{n+1}\sim P_{M}(x|\bm{x}_{1:n})+\alpha_{I}\left(P_{m}(x|\bm{x}_{1:n})-P_{M}(x|\bm{x}_{1:n})\right)$.
10:   else
11:    Compute sample space ${\mathcal{U}}_{n}$ as in Eq. (5), and sample $x_{n+1}\sim P_{M}(x|\bm{x}_{1:n})+\alpha_{U}\left(P_{m}(x|\bm{x}_{1:n})-P_{M}(x|\bm{x}_{1:n})\right)$.
12:   end if
13:   $n\leftarrow n+1$
14:   if $n\bmod b=0$ then
15:    Compute match ratio $\beta_{n/b}$, and update decoding scheme and strength parameters using Algorithm˜2.
16:   end if
17:   if $x_{n}\neq\tilde{x}_{t}$ then
18:    Exit for-loop.
19:   end if
20:  end for
21:end while
5 Experiment
This section evaluates the effectiveness, helpfulness and efficiency of the proposed method SSD. An implementation of SSD has been made available at https://github.com/k-k1w-w1x-x/Speculative-Safety-Aware-Decoding.
5.1 Experiment Setup
Models.
We evaluate SSD using open-source large models, namely, Vicuna-7b Chiang et al. (2023), Llama2-7b-chat and Llama2-13b-chat Touvron et al. (2023), as target models, and a small model TinyLlama-1.1B-Chat (Zhang et al., 2024a) as the expert model. Notice that our method only fine-tunes TinyLlama to possess the deep safety alignment property, following the method in Qi et al. (2025). Appendix C shows similar experimental results on more recent Llama3 models (Dubey et al., 2024).
Baselines.
Since we aim to strength an LLM with the additional safety property, i.e., deep safety alignment in this paper, we first test our method on the Harmful HEx-PHI dataset proposed by Qi et al. (2025). Specifically, we evaluate SSD with harmful prefixes of 10, 20, and 40 tokens. In addition, we assess the robustness of our method against other jailbreak attacks, including GCG (Zou et al., 2023), DeepInception (Li et al., 2023), and PAIR (Chao et al., 2023). To evaluate a defense method when an attacker submits harmful prompts directly to LLMs, we also consider two malicious-query benchmark datasets: Advbench (Zou et al., 2023) and HEx-PHI (Qi et al., 2023). Regarding defense methods, we compare SSD with SafeDecoding (Xu et al., 2024) and the original aligned model for prefilling attacks. We also fine-tune the original LLM with the augmented harmful dataset (Qi et al., 2025), denoted as Deep-Align in this paper. This further fine-tuned LLM serves as an oracle baseline. For other datasets, we consider three defense methods: Paraphrase (Jain et al., 2023), ICD (Wei et al., 2023b), and Self-Exam (Helbling et al., 2023).
Table 1: ASR (%) and harmful score for different models under prefilling attack.
[…]
  10 tokens   20 tokens   40 tokens
Llama2-7b No Defense 33.03% (3.13) 34.24% (3.40) 34.55% (3.44)
SafeDecoding 33.94% (3.21) 34.55% (3.47) 33.33% (3.49)
Deep-Align 1.20% (1.14) 4.50% (1.28) 10.00% (1.54)
SSD 3.64% (1.48) 5.76% (1.56) 9.70% (1.85)
Llama2-13b No Defense 25.15% (2.69) 32.42% (3.15) 27.88% (3.18)
SafeDecoding 25.15% (2.77) 30.61% (3.18) 30.30% (3.19)
SSD 3.33% (1.38) 5.45% (1.53) 8.18% (1.81)
Vicuna No Defense 68.18% (4.50) 68.79% (4.54) 65.76% (4.41)
SafeDecoding 64.55% (4.42) 68.79% (4.48) 65.15% (4.38)
Deep-Align 20.61% (2.50) 26.67% (2.84) 25.45% (2.97)
SSD 10.91% (1.88) 10.30% (1.80) 14.55% (2.05)
Evaluation Metrics.
For safety evaluation, we adopt the evaluation pipeline from Qi et al. (2023), utilizing Qwen-max (Bai et al., 2023) as a judge to automatically assess the safety of the generated outputs. The judge LLM assigns scores to the responses based on both the questions and the content of the answers, with scores ranging from 1 to 5. Here 1 indicates completely harmless and 5 indicates highly harmful. For each safety evaluation benchmark, we report the average Attack Success Rate (ASR) and the average harmful score.
For utility, we consider two datasets: GSM8K (Cobbe et al., 2021) and Just-Eval (Lin et al., 2023). GSM8K is designed to test the model’s ability to solve complex mathematical problems. Notice that TinyLlama demonstrates very low accuracy on GSM8K and this task would challenging for the composite model. We use standard ROUGE-1 score for the accuracy on GSM8K, in line with existing evaluation practices. Similar to Xu et al. (2024), we use 1,000 diverse instructions from Just-Eval to evaluate the outputs in terms of helpfulness, clarity, factuality, depth, and engagement.
A critical challenge for safety mechanisms is to avoid false positives or over-refusal responses, where the model refuses to answer legitimate queries that may touch on sensitive but harmless topics (Cui et al., 2024; Röttger et al., 2023). To quantify this, we evaluate our method on Llama2-7b using the XSTest benchmark (Röttger et al., 2023). We report the False Refusal Rate (FRR), the percentage of prompts that the model incorrectly refuses to answer, under two evaluation settings: using an LLM as a judge and via string matching. A lower FRR indicates a better ability to distinguish between harmful and merely sensitive contents, thus preserving the model’s utility on a wider range of topics.
To assess inference efficiency, we use the Average Token Generation Time Ratio (ATGR) metric:
[…]
ATGR considers the varying token lengths produced by defense methods. For our experiments, we sample 10 prompts from each of the datasets: Harmful HEx-PHI, GCG, PAIR, Just-Eval, and GSM8K, to simulate diverse real-world scenarios.
SSD Setting.
For Llama2-7b and Llama2-13b, we set the hyperparameters $\alpha_{I}=0.3$ and $\alpha_{U}=0.8$ for the $\mathrm{Intersection}$ and $\mathrm{Union}$ schemes, respectively. For Vicuna, we set $\alpha_{I}=0.45$ and $\alpha_{U}=2$, as we observe that Vicuna exhibits poorer defense capabilities than Llama2 models. We employ greedy sampling as the decoding strategy and adopt the algorithm in Stern et al. (2018) as the speculative sampling method. Due to space limit, we leave other parameter choices in Appendix˜A.
5.2 Experimental Results
Table 2: Just-Eval and GSM8K scores of different defense methods on three LLMs.
Model Defense Just-Eval (1–5) $\uparrow$ GSM8K (%) $\uparrow$
Helpfulness Clarity Factuality Depth Engaging
Llama2-7b No Defense 4.15 4.79 4.52 4.04 4.55 15.9
SafeDecoding 3.87 4.74 4.41 3.88 4.38 14.6
Deep-Align 4.00 4.74 4.44 3.91 4.47 11.6
Ours 4.08 4.78 4.44 3.98 4.53 13.7
Llama2-13b No Defense 4.40 4.86 4.62 4.23 4.71 32.3
SafeDecoding 4.16 4.81 4.54 4.16 4.55 31.5
Ours 4.36 4.88 4.62 4.20 4.69 26.5
Vicuna No Defense 4.12 4.60 4.29 3.69 3.93 24.6
SafeDecoding 3.95 4.69 4.42 3.46 3.88 15.5
Deep-Align 3.95 4.70 4.41 3.80 4.36* 11.0
Ours 3.98 4.62 4.28 3.55 3.87 22.44
SSD Transfers the Deep Safety Alignment Property. As shown in Table˜1, our method SSD consistently achieves stronger robustness to prefilling attacks than the original model and the method that lacks deep alignment property (i.e., SafeDecoding). Furthermore, when benchmarked against Deep-Align (which directly fine-tunes the large model with augmented harmful dataset), SSD achieves a close performance on Llama2-7b and even outperforms Deep-Align on Vicuna. These findings demonstrate that SSD successfully transfers the deep safety alignment property to the output responses. Besides, for less secure models like Vicuna, SSD can achieve a better safety alignment performance than directly fine-tuning the original model with supervised alignment objective.
SSD Maintains Utility.
Table 2 presents the Just-Eval and GSM8K scores on the three LLMs. We observe that the utility of SSD is largely intact. For Just-Eval, SSD incurs less than 4% decreases w.r.t. all the dimensions, compared to the original model. Most of the degradations are within 2%. Regarding GSM8K, we observe that Deep-Align substantially degrades the model’s ability to solve complex mathematical problems, particularly for Vicuna, whereas SSD effectively preserves the original model’s capability. This indicates that fine-tuning less secure models with supervised alignment objective can degrade utility performance.
**footnotetext: In Table 2, the Engaging score of Deep-Align on Vicuna is unusually high. We believe that this behavior is mainly due to the utility dataset used by Deep-Align fine-tuning, which influences Vicuna’s conversational style. See Appendix D for a detailed analysis.
SSD Avoids Over-refusal Responses on Borderline Queries.
As shown in Table˜3, SSD demonstrates a strong ability to handle borderline cases. Compared to the No Defense baseline, SSD introduces only a minimal increase in false refusals, indicating that its safety enhancement is well-targeted. In contrast, other defense methods exhibit higher FRRs, suggesting they are prone to over-refusal behavior. This result highlights that SSD effectively defends against genuine threats while maintaining helpfulness on nuanced, sensitive topics.
Table 3: False Refusal Rate (%) on the XSTest benchmark for the Llama2-7b model. A lower rate indicates fewer incorrect refusals of harmless, sensitive prompts.
Method False Refusal Rate $\downarrow$
LLM as judge String matching
No Defense 20.00% 26.80%
Deep-Align 29.60% 49.20%
SafeDecoding 28.40% 65.60%
SSD (Ours) 20.80% 30.40%
SSD Is Efficient.
Table˜4 reports the Average Token Generation Time Ratio (ATGR), comparing the inference speed of our method, SSD, against other defense strategies like Deep-Align and SafeDecoding. The results validate that SSD can successfully accelerate decoding ($\text{ATGR}<1$), particularly for larger models. In contrast, SafeDecoding’s ATGR is greater than $1$ because its design of safety checks adds computational overhead and cannot utilize the auxiliary model’s outputs to accelerate inference. It is also worth noting that the ATGRs does not account for the cost of fine-tuning, which demands greater computational resources for large models.
Table 4: Comparison of Average Token Generation Time Ratio (ATGR) across different defense methods.
Vicuna Llama2-7b Llama2-13b
SSD $\times 0.92$ $\times 0.89$ $\times 0.71$
SafeDecoding $\times 1.07$ $\times 1.03$ $\times 1.02$
Deep-Align $\times 1.00$ $\times 1.00$ $\times 1.00$
SSD Demonstrates Robustness Against Other Jailbreak Attacks.
As shown in Qi et al. (2025), deep safety alignment can also mitigate other types of jailbreak attacks to certain extent. In this experiment, we evaluate this ability of SSD against the following attack methods: GCG, PAIR, and DeepInception. We compare our approach with several existing defense strategies, including three non-decoding-time methods: Paraphrase, ICD, and Self-Examination. Due to space limit, we report the detailed results in Table˜6 in Section˜B.2.
On both Llama2-7b and Llama2-13b, the defense performance of SSD matches or surpasses the strongest baselines under three jailbreak attacks and two harmful datasets, demonstrating its robustness on the Llama2 family of models. On Vicuna, SSD achieves 0% ASR on the harmful datasets and outperforms Paraphrase and ICD across the three jailbreak attacks, yet its defense performance is slightly worse than DeepAlign and SafeDecoding. A potential reason is that TinyLlama and Vicuna differ substantially in conversational style. How to further handle this difference is beyond the scope of the present work and is left as a future work.
5.3 Ablation Study
We conduct an ablation analysis on the strength hyperparameters $\alpha_{I}$ and $\alpha_{U}$ used in SSD, as shown in Table˜5. Among them, $\alpha_{I}=0.3$ and $\alpha_{U}=0.8$ are our default parameter choices. Here the target model is Llama-7b and the harmful prefix length for the prefilling attack is 20 tokens. The results show that the safety of SSD remains at a high level for $\alpha_{I}\in[0.3,0.6]$. As $\alpha_{I}$ increases, the ASR of prefilling attack slightly increases. This can be attributed to that with an increasing $\alpha_{I}$, the match ratio may also be higher for certain harmful questions and affect the defense performance. Regarding $\alpha_{U}$, we find that its effect on the ASRs of all three attack methods remains negligible when $\alpha_{U}>0.8$.
6 Concluding Remarks
In this work, we study the problem of strengthening an existing LLM with new safety alignment properties, without tuning the model’s parameters. We propose SSD, a lightweight and efficient decoding-time approach, which employs match ratio to quantify jailbreak risks and to dynamically switch between decoding schemes to prioritize either utility or safety. Experimental results show that SSD successfully strengths the model with the desired safety property, while being helpful and efficient. A future direction is to utilize more powerful strategies of speculative sampling, e.g., Miao et al. (2024); Li et al. (2024), to further improve the flexibility and efficiency of our approach.
Table 5: SSD with different strength parameters $\alpha_{I},\alpha_{U}$.
[…]
0.3 0.6 6.36% (1.62) 8% (1.60) 4% (1.30)
0.3 1.5 5.45% (1.48) 6% (1.34) 4% (1.26)
0.3 2.0 5.15% (1.48) 6% (1.34) 4% (1.26)
0.3 0.8 4.50% (1.28) 6% (1.44) 4% (1.30)
0.4 0.8 5.45% (1.52) 4% (1.36) 4% (1.28)
0.5 0.8 7.58% (1.54) 4% (1.28) 4% (1.24)
0.6 0.8 8.18% (1.58) 2% (1.20) 2% (1.16)
7 Limitations
In this paper, the largest model on which we evaluate the proposed method is Llama2-13b. Due to limited GPU resources, we cannot conduct experiments to validate how our method applies to larger models like Llama2-70B. Another limitation of the current work is on the difference of conversational style between the original and the expert models, e.g., Vicuna and TinyLlama, which we do not investigate further. Future research can take into account this difference of conversational style to improve performance w.r.t. both model utility and safety for more practical scenarios.
8 Ethics Impact
This paper develops a lightweight and efficient method to strengthen existing LLMs with additional safety properties to defend against new jailbreak attacks. We empirically show that the developed method can effectively and efficiently equip the LLM with the deep safety alignment property and further improve the original model’s capacity against many types of jailbreaks. This research aims to enhance the safety of large models and to contribute positively to the broader field of AI research. We remark that the development of SSD only uses publicly available jailbreak prompts and do not create new ones. In the paper, only one jailbreak input query is exhibited in an abstracted way, for illustration purpose. We have also released our code to facilitate red-teaming efforts on LLMs.
Acknowledgment
This work is supported by the Innovation Project of Institute of Computing Technology, Chinese Academy of Sciences under Grant No. E561130 and the Strategic Priority Research Program of the Chinese Academy of Sciences under Grant No. XDB0680101.
Appendix A Decoding Scheme and Parameter Update
For every $b$ output tokens, Algorithm˜2 computes the running match ratio $\beta_{n/b}$ and determines the decoding scheme for next $b$ tokens. As seen from Figure˜3, the two models behave more similarly when conditioned on more output tokens. As such, if the decoding scheme keep unchanged, we adjust the match ratio threshold, and also the strength parameter in the $\mathrm{Intersection}$ state, in an annealing way, to maintain model utility. If the decoding scheme changes, these parameter values are reset.
Algorithm 2 Parameter Update
0: current match ratio $\beta_{n/b}$, initial match ratio threshold $\beta^{0}$, minimum strength weight $\alpha_{I}^{\min}>0$, decay parameters $\beta^{d},\alpha^{d}>0$
1:if $n/b=1$ then
2:  $\beta_{th}\leftarrow\beta^{0}$
3:end if
4:if $\beta_{n/b}\leq\beta_{th}$ then
5:  Set decoding scheme to $\mathrm{Intersection}$.
6:else
7:  Set decoding scheme to $\mathrm{Union}$.
8:end if
9:if decoding scheme unchanged then
10:  $\beta_{th}\leftarrow\max\left(0,\beta_{th}-\beta^{d}\right)$
11:  if $\mathrm{Intersection}$ then
12:   $\alpha_{I}\leftarrow\max\left(\alpha_{I}^{\min},\;\alpha_{I}-\alpha^{d}\right)$
13:  end if
14:else
15:  $\beta_{th}\leftarrow\beta^{0}$
16:  $\alpha_{I}\leftarrow\alpha_{I}^{0}$
17:end if
Appendix B Detailed Experimental Setups and Results
B.1 Other Parameter Choices
In our experiments, the choices of the parameters in Algorithm˜2 are: $\beta^{0}=0.6,\beta^{d}=0.1,\alpha_{I}^{{\min}}=0.3,\alpha^{d}=0.15.$. In addition, we set target sample space size to $c=10$. The lookahead in speculative sampling is $T=3$ and the bin size for computing the match ratio is $b=7$.
B.2 Experimental Results with Other Types of Jailbreak Attacks
As shown in Table 6, SSD demonstrates good defense effectiveness against all five types of jailbreak attacks on the Llama2 series models. On Vicuna, although it achieves almost 0% ASR on HEx-PHI and Advbench, and performs better than several existing defense methods on the other three types of jailbreak attacks, there is still a slight performance gap compared to Deep-Align and SafeDecoding.
Table 6: ASR (%) and harmful score for different defense methods on five benchmark datasets.
[…]
Llama2-7b No Defense 4% (1.34) 20% (2.40) 4% (1.22) 0% (1.02) 0% (1.00)
SafeDecoding 4% (1.18) 0% (1.00) 0% (1.00) 0% (1.04) 0% (1.00)
Paraphrase 0% (1.10) 2% (1.10) 0% (1.02) 0% (1.05) 0% (1.02)
ICD 0% (1.00) 4% (1.16) 0% (1.00) 0% (1.04) 0% (1.00)
Self-Examination 0% (1.00) 8% (1.40) 2% (1.08) 0% (1.00) 0% (1.00)
Deep-Align 4% (1.18) 2% (1.14) 0% (1.02) 0% (1.00) 0% (1.00)
SSD 4% (1.30) 6% (1.44) 0% (1.10) 0% (1.00) 0% (1.00)
Llama2-13b No Defense 4% (1.40) 2% (1.08) 0% (1.16) 0% (1.06) 0% (1.00)
SafeDecoding 4% (1.26) 0% (1.02) 0% (1.14) 0.30% (1.06) 0% (1.00)
Paraphrase 0% (1.10) 0% (1.08) 0% (1.24) 0% (1.05) 0% (1.00)
ICD 0% (1.00) 0% (1.02) 0% (1.00) 0% (1.03) 0% (1.00)
Self-Examination 2% (1.22) 0% (1.00) 0% (1.16) 0% (1.05) 0% (1.00)
SSD 2% (1.26) 0% (1.06) 0% (1.16) 0% (1.04) 0% (1.00)
Vicuna No Defense 80% (4.72) 86% (4.86) 58% (4.34) 10.30% (1.72) 4% (1.38)
SafeDecoding 4% (1.28) 0% (1.14) 0% (1.08) 0.91% (1.17) 0% (1.00)
Paraphrase 24% (2.32) 42% (3.18) 40% (3.92) 13.33% (1.94) 2% (1.22)
ICD 24% (2.34) 48% (3.16) 38% (4.18) 3.94% (1.23) 0% (1.00)
Self-Examination 8% (1.66) 6% (1.48) 48% (4.04) 7.27% (1.57) 0% (1.22)
Deep-Align 2% (1.08) 0% (1.00) 0% (1.02) 0% (1.01) 0% (1.00)
SSD 18% (2.14) 10% (1.40) 8% (2.30) 0.91% (1.18) 0% (1.00)
B.3 Setup of Attack Methods
We utiliz Harmful HEx-PHI dataset (Qi et al., 2025) to conduct the prefilling attack. This dataset consists of 330 harmful instructions extracted from the HEx-PHI safety benchmark, with harmful answers generated using a jailbroken version of GPT-3.5-Turbo. In this study, the dataset is used for prefilling attacks, following the same methodology as Qi et al. (2025) by concatenating harmful answers of varying lengths (i.e., 10, 20, and 40 tokens) with the harmful instructions. For GCG Zou et al. (2023) and PAIR Chao et al. (2023), we follow Chao et al. (2023); Xu et al. (2024) and utilize 50 distinct representative harmful queries [^0] https://github.com/patrickrchao/JailbreakingLLMs from Advbench Zou et al. (2023) to generate specific attack prompts for each model. Due to limitation on computational resources, we use top-$k=64$ setting for the GCG attack on Llama2-13b-chat. The rest hyperparameters are consistent with those described in the original paper. For DeepInception, we apply the ready-to-use template prompt provided in the Github repository [^1] https://github.com/tmlr-group/DeepInception. HEx-PHI safety benchmark contains 330 harmful instructions specifically designed for LLM harmfulness evaluation.
B.4 Setup of Deep-Align
We train the Llama2-7b and Vicuna models with deep safety alignment properties using the same hyperparameters and datasets as the default settings in Qi et al. (2025). For TinyLlama used as the small expert model $m$, we adjust the learning rate of the hyperparameters to $2\times 10^{-4}$.
Appendix C Experimental Results on Llama3 Models
We conduct experiments on a more recent LLM Llama3.1-8b-Instruct (Dubey et al., 2024), to further validate the effectiveness of SSD. Here we use Llama3.2-1b-Instruct as the expert model because it shares the same vocabulary with Llama3.1-8b-Instruct. The results regarding safety and utility are presented in Table˜7 and Table˜8, respectively.
As shown in Table˜7, the original Llama3.1-8b model is highly vulnerable to prefilling attacks, with ASR of approximately 50%. Deep-Align provides a strong defense performance while the safety performance of SSD is close. This confirms that SSD effectively transfers the deep safety alignment property to this newer model. However, the utility evaluation in Table˜8 highlights the critical trade-offs between safety and utility. Our method shows much better utility performance compared with Deep-Align, particularly for the GSM8K score. Overall, SSD offers a comparable performance of the desired safety property while ensuring that the model’s utility remains largely intact.
Table 7: ASR (%) and harmful score for the Llama3.1-8b model under prefilling attack.
[…]
  10 tokens   20 tokens   40 tokens
Llama3.1-8b No Defense 46.73% (3.71) 49.84% (3.97) 50.16% (4.07)
Deep-Align 0.91% (1.08) 1.52% (1.14) 2.42% (1.16)
SSD 2.52% (1.59) 5.85% (1.69) 4.67% (1.71)
Table 8: Just-Eval and GSM8K scores of different defense methods on the Llama3.1-8b model.
Model Defense Just-Eval (1–5) $\uparrow$ GSM8K (%) $\uparrow$
Helpfulness Clarity Factuality Depth Engaging
Llama3.1-8b No Defense 4.23 4.89 4.80 3.97 4.35 81.8
Deep-Align 3.44 4.50 3.77 3.32 4.06 10.3
Ours 3.93 4.73 4.59 3.71 4.06 75.2
Appendix D Analysis of the Anomaly Engaging Score
In our main results in Table 2, we note that Deep-Align on Vicuna exhibits an anomalously high Engaging score of 4.36. We believe that the primary reason is that the anchor utility dataset used in Deep-Align promotes a more polite and conversational tone in Vicuna’s responses.
Table 9 provides a qualitative analysis comparing responses from both the original model (Baseline) and the Deep-Align model. Deep-Algin tends to introduce engaging openers like “Of course! I’m happy to help” or “Thank you for your question!”. To further validate this, we fine-tune Vicuna only on this utility dataset, which results in an even higher Engaging score of 4.46.
Another reason is the response length: the original model has an average response length of 179 words with Engaging score of 3.93, while the average length of Deep-Align is 227 words in Table 2. We then conduct a controlled-length experiment by constraining the output length via Prompting, and the results are reported in Table 10. We find that even with average length of 180 words, the Engaging score of Deep-Align is still much higher, suggesting that the response length is only a minor issue to the high Engaining score reported in Table 2.
Table 9: Comparison of response styles between the baseline Vicuna model and the Deep-Align model.
Query 1: What are some famous actors that started on Broadway?
Baseline There are many famous actors who got their start on Broadway. Here are a few examples…
Deep-Align Of course! I’m happy to help. Some famous actors who started their careers on Broadway include…
Query 2: How did US states get their names?
Baseline The United States is made up of 50 states…
Deep-Align Thank you for your question! The United States is a vast country…
Query 3: What is some cool music from the 1920s?
Baseline The 1920s was a decade of significant artistic and cultural change…
Deep-Align Of course! The 1920s were a time of great musical innovation…
Table 10: Engaging score of Deep-Align on Vicuna with different output length constraints. The undefended baseline model has an average length of 179 words and an Engaging score of 3.93.
Average Output Length 153 159 180 227
Engaging Score 4.24 4.29 4.31 4.42